# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 18 Mar 2026 v2

@@ Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes. We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, entirely bypassing the primorial-prime hypothesis. We further establish the sharp order
@@
and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$. We have not found this full-support statement in the literature. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma and is moreover continuous and purely singular. We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1. As a simple conservative explicit bound, for any prime $q>2^{30}$ the expected overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@ 1 Introduction
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 8. Hence the decay to zero is completely quantified and extremely slow. The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$, well within practical ranges. The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
3. Limiting distribution (Theorems 6 and 7, Corollary 2). The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular. If $G$ denotes the distribution function, then
@@
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$. The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4]. Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p. 263]; we include a self-contained proof for the reader’s convenience. Our new contributions on the distributional side are twofold. First, we establish that the support of the limiting law is exactly $0,\tfrac{1}{2}]$ by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Second, we transfer the classical shifted-prime law for $\varphi(p-1)/(p-1)$ to the prim-lwe density $c(p)$ via a transport lemma (Lemma 5).
4. Explicit bounds for cryptographic moduli (Theorem 8, Table 1). We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure. The following consequence singles out the cryptographic implication.
@@
Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then, in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ satisfies
@@
For each fixed dimension $n$, the exact expected overhead is $1/c_{n}(q)\leq 1/c(q)$.
@@
Section 2 records the precise definitions and the open question from [1]. Section 3 gives the unconditional proof that $\inf_{p}c(p)=0$. Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$. Section 5 develops the limiting distribution theory. Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$. Section 7 discusses the implications for prim-lwe.
@@
We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order. All limits are taken as $x\to\infty$ unless stated otherwise. We use the standard Vinogradov notation $A\ll B$ and $A\gg B$ synonymously with $A=O(B)$ and $B=O(A)$, respectively. $A\asymp B$ means both $A\ll B$ and $A\gg B$. We write $\operatorname{Li}(x)=\int_{2}^{x}\mathrm{d}t/\!\log t$ for the logarithmic integral, and throughout $\log$ denotes the natural logarithm. We use the notation $A=O^{*}(B)$ to mean $|A|\leq B$, i.e. an implied constant of $1$.
@@ 2 Background and the Open Question
Recall that one may equivalently bundle $n$ standard LWE secret vectors into a matrix and view decision-LWE in matrix form: given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
@@
from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7]. Via the identification $\mathbb{Z}_{q}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
@@ 3 Unconditional Proof that $\inf_{p}c(p)=0$
We also record the following simple observation.
@@ 4 Quantitative Refinements
Let
$N_{k}=\prod_{i=1}^{k}p_{i},\qquad\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x\\ p\text{ prime}\end{subarray}}\log p.$
Then
$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}\qquad(k\to\infty).$
Consequently, if
$m(x):=\max\{k\geq 1:N_{k}\leq x\},$
then
$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x\qquad(x\to\infty).$
@@ 5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp. 257–258] and Deshouillers–Hassani [4].
Our contributions in this section are:
  1. (i)
    The full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4).
  2. (ii)
    The transport of the classical shifted-prime law to the prim-lwe density $c(p)$ via Lemma 5, establishing that $c(p)$ shares the same limiting distribution as $\varphi(p-1)/(p-1)$.
We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9.
@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
Let $h$ be the strongly additive function $h(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
$0<h(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad h(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
Consequently the third Erdős–Wintner series $\sum_{|h(\ell)|>1}1/\ell$ is vacuous, while $\sum_{\ell}h(\ell)/\ell$ and $\sum_{\ell}h(\ell)^{2}/\ell$ both converge. Since $h(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied.
The same conclusions hold for $g(n)=\varphi(n)/n$ via $\log g(\ell)=-h(\ell)$.
@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
The following transport lemma is the mechanism for contribution (ii) listed at the beginning of this section.
@@
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p. 258]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [3, formula (3), p. 258], the characteristic function of $\mu_{f}$ is given by
@@
Let $y\geq 3$, and for each prime $p$ define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell$. Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. Then the set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
@@
Let $f(n):=\log(n/\varphi(n))$, and for each prime $p$ and parameter $y\geq 3$ define
@@
where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$. In particular, $\eta(y)\to 0$ as $y\to\infty$.
@@
Let
$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
@@
By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp. 257–258] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
@@
Let
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then the empirical measures
@@
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp. 257–258] for the general shift including p1p-1) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
@@
Set
$f(n):=\log\frac{n}{\varphi(n)}=\sum_{\ell\mid n}\log\frac{\ell}{\ell-1}.$
Then $f$ is strongly additive and
@@
$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)=\left(\frac{1}{2};\frac{1}{2}\right)_{\infty}\approx 0.288788<\frac{1}{2}.$

### 30 Mar 2026 v3

@@ Abstract
The prim-lwe problem, introduced by Sehrawat, Yeo, and Desmedt (Theoretical Computer Science, 886 (2021)), is a variant of the Learning with Errors problem in which the secret matrix is required to have a primitive-root determinant. The dimension-uniform reduction constant is $c(p)=\inf_{n\geq 1}c_{n}(p)$, where $c_{n}(p)$ is the exact density of $n\times n$ matrices over $\mathbb{F}_{p}$ with primitive-root determinant. Sehrawat, Yeo, and Desmedt asked whether $\inf_{p\text{ prime}}c(p)=0$, observing that an affirmative answer would follow from the conjectural infinitude of primorial primes. We resolve this question unconditionally using only Dirichlet’s theorem and Mertens’ product formula, bypassing the primorial-prime hypothesis. We further establish the sharp order
@@
and show that the limiting distribution of $c(p)$ over the primes has support exactly $[0,1/2]$. This full-support property appears to be new. The law coincides with the classical shifted-prime distribution of $\varphi(p-1)/(p-1)$ via a transport lemma, is moreover continuous and purely singular, and the limiting measure has Hausdorff dimension 0. The moments of the limiting distribution are given by convergent Euler products that uniquely determine it, and the associated Mellin transform extends analytically to the right half-plane. We also derive explicit lower bounds on $c(q)$ for primes of cryptographic interest, parameterized solely by the number of distinct prime factors of q1q-1. As a simple conservative explicit bound, for any prime $q>2^{30}$ the dimension-uniform expected overhead $1/c(q)$ is at most $1.79\log q$. On the other hand, our results show that the worst-case overhead among primes $p\leq x$ is of order $\Theta(\log\log x)$, and in particular $1/c(q)=O(\log\log q)$ pointwise.
@@ 1 Introduction
The matching lower bound follows from two observations: if $p\leq x$ is an odd prime, then $\omega(p-1)\leq m(x):=\max\{k:N_{k}\leq x\}$, and Mertens’ theorem converts this into a $1/\!\log\log x$ lower bound for $\varphi(p-1)/(p-1)$; the additional factor $\prod_{j\geq 1}(1-p^{-j})\geq P_{3}$ is handled by Theorem 11. Hence the decay to zero is quantified precisely. The implied constant is $\approx 0.31$, so primes $p$ whose predecessor p1p-1 is divisible by many small primes can still have $c(p)<0.1$. The explicit bounds in Section 6 show that standard cryptographic moduli with controlled factorization incur only a small constant-factor overhead; see Remark 1 for effectivity.
3. Limiting distribution (Theorems 6 and 7, Corollary 2, Proposition 4). The quantity $c(p)$ possesses a continuous limiting distribution over the primes whose support is exactly $0,\tfrac{1}{2}]$; the law is moreover purely singular, and Proposition 4 shows that the limiting measure has Hausdorff dimension 0. If $G$ denotes the distribution function, then
@@
the set of primes with $c(p)\leq\tau$ has relative density $G(\tau)$, and $G(\tau)\to 0$ as $\tau\downarrow 0$. The existence and continuity of $G$ follow from the shifted-prime Erdős–Wintner framework for additive functions evaluated at p1p-1 [2]; for the relevant classical background see Tenenbaum [3] and Deshouillers–Hassani [4]. Pure singularity is classical in the integer case (Erdős [5]) and carries over to shifted primes by the same method, as noted by Tenenbaum [3, p. 263]; we include a self-contained proof for the reader’s convenience. Our new contribution on the distributional side is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$, established by combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Since the Euler product $P(p)\to 1$, the prim-lwe density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ shares the same limiting distribution (Lemma 5). We also derive the moments of $G$ as convergent Euler products (Theorem 8) and establish the analytic continuation of the Mellin transform $\mathbb{E}[X^{s}]$ to the right half-plane (Theorem 10). At the right endpoint, $1-G(\tfrac{1}{2}-\varepsilon)\sim\kappa/\!\log(1/\varepsilon)$ as $\varepsilon\downarrow 0$ for an explicit constant $\kappa$ (Proposition 7), so $G$ is not Hölder of any positive order at $1/2$.
4. Explicit bounds for cryptographic moduli (Theorem 11, Table 1). We give unconditional lower bounds on $c(q)$ for primes $q$ parameterized by $\omega(q-1)$, the number of distinct prime factors of q1q-1, yielding concrete bounds for the NIST-standardized primes and other moduli whose predecessors have controlled factorization structure. The following consequence singles out the cryptographic implication: in the reduction from decision-LWE to decision-prim-lwe over $\mathbb{F}_{q}$, the dimension-uniform expected rejection-sampling overhead $1/c(q)$ admits an explicit upper bound depending only on $\omega(q-1)$.
@@
Let $q\geq 3$ be a prime modulus with $\omega(q-1)\leq K$. Then
@@
Section 2 records the precise definitions and the open question from [1]. Section 3 gives the unconditional proof that $\inf_{p}c(p)=0$. Section 4 proves the upper bound $\min_{p\leq x}c(p)\ll 1/\log\log x$. Section 5 develops the limiting distribution theory, including the Hausdorff-dimension-zero result and the moment Euler product (Section 5.3). Section 6 derives explicit lower bounds for primes of cryptographic interest and, in particular, supplies the matching lower bound $\min_{p\leq x}c(p)\gg 1/\log\log x$. Section 7 discusses the implications for prim-lwe.
@@
We write $\omega(n)$ for the number of distinct prime factors of an integer $n$, $\pi(x)$ for the number of primes up to $x$, and adopt the convention that $p_{1}=2<p_{2}=3<p_{3}=5<\cdots$ denotes the sequence of all primes in increasing order. For a prime $p$, we write
$P(p):=\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{p^{j}}\Bigr)$
for the matrix-count Euler product; thus $c(p)=\frac{\varphi(p-1)}{p-1}\,P(p)$. All limits are taken as $x\to\infty$ unless stated otherwise. We use the standard Vinogradov notation $A\ll B$ and $A\gg B$ synonymously with $A=O(B)$ and $B=O(A)$, respectively. $A\asymp B$ means both $A\ll B$ and $A\gg B$. We write $\operatorname{Li}(x)=\int_{2}^{x}\mathrm{d}t/\!\log t$ for the logarithmic integral, and throughout $\log$ denotes the natural logarithm. We use the notation $A=O^{*}(B)$ to mean $|A|\leq B$, i.e. an implied constant of $1$.
@@ 2 Background and the Open Question
In the matrix form of decision-LWE, given $\mathbf{A}\in\mathbb{Z}_{q}^{w\times n}$, one must distinguish
@@
from uniform over $\mathbb{Z}_{q}^{w\times n}\times\mathbb{Z}_{q}^{w\times n}$, where $\mathbf{S}\in\mathbb{Z}_{q}^{n\times n}$ is the secret matrix and $\mathbf{E}\in\mathbb{Z}_{q}^{w\times n}$ is the error matrix [7]. Via the identification $\mathbb{Z}/q\mathbb{Z}\cong\mathbb{F}_{q}$ for prime $q$, decision-prim-lwe [1] is the variant in which the secret matrix is additionally required to lie in $M_{n}^{\mathrm{prim}}(\mathbb{F}_{q})$.
@@ 3 Unconditional Proof that $\inf_{p}c(p)=0$
We also record the following observation.
@@ 4 Quantitative Refinements
Let $N_{k}:=\prod_{i=1}^{k}p_{i}$ denote the $k$-th primorial, $\vartheta(x):=\sum_{\begin{subarray}{c}p\leq x,\;p\text{ prime}\end{subarray}}\log p$ the Chebyshev function, and $m(x):=\max\{k\geq 1:N_{k}\leq x\}$.
@@
As $k\to\infty$,
$\log N_{k}=\vartheta(p_{k})\sim p_{k},\qquad p_{k+1}\sim p_{k}.$
As $x\to\infty$,
$\log N_{m(x)}\sim\log x,\qquad p_{m(x)}\sim\log x.$
@@ 5 The Limiting Distribution of $c(p)$
We now turn to the distributional theory of $c(p)$ over the primes. The classical existence and continuity of the limiting distribution of $\varphi(p-1)/(p-1)$ over primes follow from the shifted-prime Erdős–Wintner framework; for the shift a=1a=-1 used here, direct references are Tenenbaum [3, pp. 262–263] and Deshouillers–Hassani [4].
Throughout this section we write
$f(n)\;:=\;\log\frac{n}{\varphi(n)}\;=\;\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}$ (1)
for the strongly additive function that records $\log(n/\varphi(n))$.
Our contribution in this section is the full-support conclusion: the limiting law has support exactly $0,\tfrac{1}{2}]$ (Theorem 6), combining the left non-differentiability result of [4, Theorem 1.1] with the density of $\{\varphi(m)/m:m\text{ even}\}$ in $[0,1/2]$ (Lemma 4). Since $P(p)\to 1$, the prim-lwe density $c(p)$ shares the same limiting distribution (Lemma 5). We supply a self-contained proof of the shifted-prime pure singularity in Lemma 9; in Section 5.3 we show that the limiting measure has Hausdorff dimension 0 and derive the moments of $G$ as convergent Euler products.
@@ 5.1 The distribution of $\varphi(p-1)/(p-1)$
The function $f$ defined in (1) is strongly additive, with $f(n)=\log(n/\varphi(n))=-\log(\varphi(n)/n)$. For every prime $\ell$,
$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
The three Erdős–Wintner convergence conditions are satisfied, as is the continuity criterion of Lemma 6. The same conclusions hold for $g(n)=\varphi(n)/n$.
Proof.
Since $|f(\ell)|\leq\log 2<1$ for every prime $\ell$, the series $\sum_{|f(\ell)|>1}1/\ell$ is vacuous. The estimate $f(\ell)=1/\ell+O(1/\ell^{2})$ gives $\sum_{\ell}f(\ell)/\ell<\infty$ and $\sum_{\ell}f(\ell)^{2}/\ell<\infty$. Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the continuity criterion of Lemma 6 is satisfied. The case $g(n)=\varphi(n)/n$ follows from $\log g(\ell)=-f(\ell)$. ∎
@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Since the Euler product $P(p)\to 1$, the density $c(p)=\frac{\varphi(p-1)}{p-1}P(p)$ inherits the limiting distribution of $\varphi(p-1)/(p-1)$. The following lemma records this fact.
@@
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p. 262]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [3, formula (3), p. 262], the characteristic function of $\mu_{f}$ is given by
@@
For $y\geq 3$ and each prime $p$, define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime},\;\ell\mid(p-1)\end{subarray}}\ell$, and write $P^{+}(n)$ for the largest prime factor of $n$.
@@
Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. The set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
@@
Let $f$ be as in (1), and for each prime $p$ and parameter $y\geq 3$ define
@@
where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$.
@@
Let $f$ be the strongly additive function defined in (1). The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
@@
By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp. 262–263] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
@@
Let $f$ be as in (1). Then the empirical measures
@@
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp. 262–263] for the general shift including p1p-1) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
@@
Let $f$ be as in (1). Then $f$ is strongly additive and
@@
$c(2)=\prod_{j\geq 1}\Bigl(1-2^{-j}\Bigr)\approx 0.288788<\frac{1}{2}.$